OpenStudy (anonymous):

Given the quadratic function f(x) = x² – 12x + 36, find a value of x such that f(x) = 9.

OpenStudy (helder_edwin):

it is just a quadratic equation \[ \large 9=f(x)=x^2-12x+36 \]

OpenStudy (anonymous):

is 9 the value of x?

OpenStudy (helder_edwin):

solve the equation. u didn't even try!

OpenStudy (anonymous):

x=9, x=3

OpenStudy (anonymous):

x^2-12x+36=9 x^2-12x+27=0 (x-9)(x-3)=0 x=9, x=3

OpenStudy (helder_edwin):

yes well done.

OpenStudy (anonymous):

thank you

OpenStudy (helder_edwin):

u r welcome